Frequency and Wavelengths


1.What is the frequency of green

light that has a wavelength of 5.5 x 10 -7 m?

 

2. What is the wavelength of a microwave that has a frequency of 4.2 x 10 8 Hz?

 

3. Find the green of light from no.1 if its wavelength changes to 4.2 x 10 -7 m.

 

4. Red light has a frequency of 4.0 x 10 14 Hz.

  a. Find the speed of red light in a vacuum if its wavelength is 7.5 x 10 -7 m.

  b. As the red light above enters glass its wavelength changes to 5.8 x 10 -7 , but its

frequency remains constant. What is the speed of red light in this glass?


1
Expert's answer
2021-01-23T14:08:17-0500

1)


"f=\frac{c}{f}\\f=\frac{3\cdot10^8}{5.5\cdot10^{-7}}=5.5\cdot10^{14}\ Hz"

2)


"\lambda=\frac{c}{f}\\\lambda=\frac{3\cdot10^8}{4.2\cdot10^8}=0.714\ m"

3)


"f=\frac{c}{f}\\f=\frac{3\cdot10^8}{4.2\cdot10^{-7}}=7.1\cdot10^{14}\ Hz"

4) a)


"c=f\lambda=(4\cdot10^{14})(7.5\cdot10^{-7})=3.0\cdot10^{8}\frac{m}{s}"

b)


"c=f\lambda=(4\cdot10^{14})(5.8\cdot10^{-7})=2.3\cdot10^{8}\frac{m}{s}"